To build a set of book shelves, a carpenter needs 30 identical boards, each 2 feet 7 inches long. For this type of wood, the lumber yard sells only 12 foot long boards. How many 12 foot long boards must the carpenter purchase?
A.6 B.7 C.8 D. 9 E More than 9

Answer by Theo(13342)

each board is 2 feet and 7 inches long.

this is equivalent to (2 * 12) + 7 = 31 inches long.

the lumber yard sells only 12 foot boards.

this is equivalent to (12 * 12) = 144 inches long.

he can get 144/31 = 4.64516129 shelves out of each board.

since he needs a full shelf, that means he can only get 4 shelves out of each board.

he needs 30 shelves / 4 equals 7.5 boards.

since he can't buy a partial board, he has to buy 8 boards.

8 * 4 = 32 which means he will have 2 extra shelves.

your answer is selection (C).